The time has come yet again for the annual overhaul of the SPCs. With volunteers from the Scripps diver community, the cameras were removed from the pier piling, hoisted to the pier deck, and thoroughly cleaned. The cables and underwater wipers were replaced, and then the refreshed instrument was lowered down and reinstalled. Many thanks to the hard work of our divers Natalia, Rex, Rich, and Taylor!
Early in November, the SPCs were pulled from the pier and brought back to the lab for a full overhaul. One of the internal computers had suffered thermal damage from a damaged cooling fan, so after some replacements, upgrades, and realigning, the SPCs were successfully remounted yesterday. They are once again onine!
Recently, biofouling on the ports began ot cut off the field of view, showig that it was time for another cleaning. Scripps divers were able to clean off the ports, as well as replace a cable that had ceased to function. Thanks to their hard work, the SPC is once again capturing and saving live data. How does a cable break happen? Generally it starts with a small hole in the jacket from a barnacle. All it takes is a smal gap for seawater to access one of the conductors in the cable. In this case, it was the ground line that got exposed. Once open to the water, galvanic corrosion sped along by the operating voltage then make short work of the exposed conductor.
After a maintenance dive to remove a buildup of biofouling, and extensive work on the server, both the cameras and the website are back online!
The backend of the SPC website is undergoing some sorely needed upgrades and revitalization as the servers are being updated and replaced with more secure systems. Sadly, the plankton viewer will be down until this work can be completed. The data is safe, and the viewer will be back up as soon as possible!
Pichaya and Eric removed the entire SPC frame yesterday morning to take care of some major maintenance. After only three months since the last surface cleaning session, the whole system had gotten pretty fuzzy. Check out all that growth!
Aside from giving the SPC a shave, the major reason for pulling the instrument was to reinstall the port wipers. Putting the wipers on the frame is too challenging for a diver to do underwater, so we had to wait for a good swell window to bring it to surface.
Once the SPC was on the pier, Devin and Alejandra joined in for some good ol’ barnacle smashing and gunk scraping. After the system was reasonably clean, we attached wipers. A good amount of time went into aligning the blades – each one needs to sweep across two ports with enough pressure to remove any sediment or critters.
With the wipers installed and tested, the team decided to change the u-bolts for a clamp system. The u-bolts have held the camera housings to the frame for several years. The threads were corroded from years of exposure and ran the risk of coming loose. Moreover, the clamps are a little easier to install and leave more space for cables.
After re-securing all the cables and the copper mesh cage, Pichaya and Eric returned the fully assembled system to its spot on the pier. All told, the camera was out of the water for about 6 hours. Between the cage and wipers, we hope the SPC will not need to come to the surface again for another 6 months or so.
The SPC immediately started collecting a huge number of images of diatom chains! Go to the image viewer for more of those. You can also look at our historical data and to see how the plankton population shifts over time.
Check out more of Pichaya’s pictures of our cleaning session on Scripps Pier.
Paul and Eric cleaned the cameras today. Both strobe ports were heavily fouled.
We pulled the cameras today to perform maintenance. We hope to get them back in the water as soon as possible!
The SPCP2 (Phytoplankton) camera was down for two days due to a problem with the network cable. It is now back up and running and we will hopefully recover the data from the local camera storage.
There has been a great deal of fouling due to growth of muscles on the camera in the past few months. Today Pichaya and Paul cleaned the cameras and removed large amounts of small muscles from nearly everywhere on the system. Likely both wipers had been fouled and where not operating for the 10 days. Given the extensive fouling on the cameras, they will be temporarily removed for a major service in the near future.